Assignment 1, Mark:  (10 %)

1. If 2g of NaCO3 dissolved in 1000ml of water, what is the normality of the solution?

2. How many equivalents of phosphoric acid are contained in 300 mL of 4.00 M phosphoric acid (Mwt = 98 g/mol)? (Assume the acid is to be completely neutralized by a base.)

3.                What is the pH after 20.0 mL of 0.010 M HCl is added to A buffer that contains both 0.250 M HF and 0.500 M NaF?

4.                Calculate the pH of the solution, when 50.0 ml of 0.5 M HNO2 (Ka = 4.0x10-4) mixed with 25.0 ml of 1.0 M KOH solution.

5. If 50 ml of 0.5M HCl solution titrated with 60 ml of 0.2 M of NaOH, what would be the pH of the final solution? 



1
Expert's answer
2021-05-04T14:28:18-0400

1.) Normality = Molarity "\times" n-Factor

Normality "= \dfrac{\dfrac{2}{106}}{\dfrac{1000}{1000}} \times 2 = 0.03N"


2.) "4.80eq"

3.) In this question we can use and in the other parts also.

"pH = -log[H]^+"

4.) We can use the formula,

"pH = pK_a +log \dfrac{[H^-]}{HA}"


5.) "pOH = pK_b +log \dfrac{Salt}{Base}"